In how many ways can a group of 10 people be divided into three groups consisting of 4,3, and 3 people?
Answer by mananth(16946)


We use the formula of combinations nCr where n represents the total number of items and r
is the number of items taken at the time
First group 4 people from the 10
C(10,4)=210
Second group 3 from balance (10-4)=6
C(6,3)=20
Third group 3 from 3 , 1 group
Total ways = 210 * 20 = 4,200 ways.